OpenStudy (anonymous):

Write the equation of the line with a slope of -1/3 and passing through the point (6, -4).

OpenStudy (anonymous):

You will use the point-slope formula: \[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

I don't know math what so ever

OpenStudy (anonymous):

M is the slope (-1/3). y1 is your y-value (-4) and x1 is your x-value(6).

OpenStudy (anonymous):

so that's my final answer?

OpenStudy (anonymous):

I'm here to help haha. I can't just give you the answer. It's against the rules. So I'll try to explain as best as I can.

OpenStudy (anonymous):

Lol, you can break rules

OpenStudy (anonymous):

Let's just try. This is really easy. You just plug in and solve. So, if you have your m (-1/3), your y1 (-4) and your x1 (6), you can plug in to the equation.

OpenStudy (anonymous):

You should get: y- (-4) = -1/3 (x-6)

OpenStudy (anonymous):

such a life saver <3

OpenStudy (anonymous):

So, simplify: y+4=-1/3x+2

OpenStudy (anonymous):

Lol we're not done haha.

OpenStudy (anonymous):

Lol what the heck its more ?

OpenStudy (anonymous):